# Semiclassical Canovaccio for Composite Operators

### v1

@@ 5.3 Fluctuation operators
$\xi_{\ell,\pm}(z)=\frac{H(z\pm\alpha,\,|\,m)}{\Theta(z,\,|\,m)}\,e^{\mp\,z\,Z(\alpha,\,|\,m)}\ ,$ (79)
where $H,\Theta,Z$ respectively denote the Jacobi Eta, Theta, and Zeta functions reviewed in appendix B, and $\alpha=\alpha(\Lambda_{1}(\ell))$ solves
@@
$\xi_{\ell,\pm}(z)=\frac{H(z\pm\alpha_{+},\,|\,m)H(z\pm\alpha_{-},\,|\,m)}{\Theta(z,\,|\,m)^{2}}e^{\mp z(Z(\alpha_{+},\,|\,m)+Z(\alpha_{-},\,|\,m))}\,$, (82)
@@
$\xi_{\ell,\pm}(z)=\prod_{j=1}^{\kappa}\frac{H(z\pm\alpha_{j}\,|\,m)}{\Theta(z,\,|\,m)}e^{\mp zZ(\alpha_{j}\,|\,m)},$ (87)
@@
$C_{1}=\delta E_{1}+\frac{1}{2{\cal T}}\sum_{\ell=0}^{\infty}\left(n_{\ell}\left[(N-1)\nu_{1,\ell}+\nu_{2,\ell}-N\nu_{0,\ell}\right]+2q_{1,\ell}\nu_{1,\ell}+2q_{2,\ell}\nu_{2,\ell}\right)\ ,$ (89)
@@
$C_{1}=\delta E_{1}+\frac{1}{2{\cal T}}\sum_{\ell=0}^{\infty}\left(n_{\ell}\left[(N-1)\nu_{1,\ell}+\nu_{2,\ell}\right]+2q_{1,\ell}\nu_{1,\ell}+2q_{2,\ell}\nu_{2,\ell}\right)\ ,$ (93)
@@
We now comment on the zero modes of the theory. It is shown that when a symmetry of the theory is broken by a classical field configuration, a given pair of stability angles vanishes dhn . In our case, the stability angle $\nu_{2,0}$ corresponds to the zero mode related to the time translation symmetry on the cylinder. The unnormalized eigenfunction is $\dot{v}(t)$ multiplied by the $\ell=0$ spherical harmonic. Indeed, one has $\Lambda_{2}(\ell=0)=1+4m$, which coincides with one of the band edges of the $L_{2}$ Lamé operator $L_{2}$. To evaluate the associated stability angle, we use that for $\ell=0$ and $d=4$ the equations (83) reduce to
@@ 5.4 Leading quantum correction: $C_{1}$
$\frac{1}{2}\sum_{\ell=0}^{\infty}n_{\ell}\nu_{2,\ell}=2{\cal T}+\frac{1}{2}\sum_{\ell=2}^{\infty}n_{\ell}\nu_{2,\ell}\ ,$ (100)
@@
$\frac{1}{2}\sum_{k=1}^{5}c_{k}\zeta(k-d)$ $=\pi-\frac{3((m-1)(3m-2)\mathbb{K}(m)+(4m-2)\mathbb{E}(m))}{(1-2m)^{3/2}\epsilon}-\frac{2(m-2)\mathbb{K}(m)+6\mathbb{E}(m)}{\sqrt{1-2m}}\ .$ (104)
@@
$-\frac{1}{2}\sum_{k=1}^{5}c_{k}|_{d=4}=\frac{(m(13-31m)+2)\mathbb{K}(m)+4\pi(1-2m)^{3/2}-(18-36m)\mathbb{E}(m)}{(1-2m)^{3/2}}\ .$ (105)
@@
$-\frac{((29-5m)m-14)\mathbb{K}(m)-5\pi(1-2m)^{3/2}+(24-48m)\mathbb{E}(m)}{(1-2m)^{3/2}}\ ,$ (106)
@@
By comparing the sum of Eq. (5.4) and Eq. (5.4) (multiplied by N1N-1) to Eq. (65), one can see that the divergent $1/\epsilon$ pole cancels in the renormalized combination $\mathcal{S}_{\rm cl}-\frac{1}{2}\sum_{\ell=0}^{\infty}n_{\ell}\left(\nu_{2,\ell}+(N-1)\nu_{1,\ell}\right)$. Dividing by the period $\cal T$} and taking into account our results for $\delta E_{1}$ in Eq. (68), we finally arrive at
$C_{1}$ $=-\frac{(1-m)m(N+8)}{8(1-2m)^{2}}+\frac{1}{8(1-2m)^{2}\mathbb{K}(m)}\Bigg[\mathbb{K}(m)(m(-7mN+26m+3N-70)+28)$
@@ 5.5 Perturbative semiclassics
$\frac{1}{2}\sum_{\ell=1}^{\infty}\sigma_{1}(\ell)$ $=\frac{11\pi m^{2}}{64}+\mathcal{O}(m^{3})=11\pi\left(\frac{\lambda n}{16\pi^{2}}\right)^{2}+\mathcal{O}\left(\left(\frac{\lambda n}{16\pi^{2}}\right)^{3}\right)\,$, (111)
$\frac{1}{2}\sum_{\ell=2}^{\infty}\sigma_{2}(\ell)$ $=\frac{51\pi m^{2}}{64}+\mathcal{O}(m^{3})=51\pi\left(\frac{\lambda n}{16\pi^{2}}\right)^{2}+\mathcal{O}\left(\left(\frac{\lambda n}{16\pi^{2}}\right)^{3}\right)\ ,$ (112)
@@
$\begin{split}\frac{\nu_{1,\ell}}{\mathcal{T}}\bigg|_{d=4}&=\ell-\frac{2(3\ell+1)}{\ell+1}\left(\frac{\lambda n}{16\pi^{2}}\right)\\ &+\frac{3\ell(\ell+2)(\ell(\ell(17\ell+43)+35)+5)-4}{\ell(\ell+1)^{3}(\ell+2)}\left(\frac{\lambda n}{16\pi^{2}}\right)^{2}+\mathcal{O}\left(\left(\frac{\lambda n}{16\pi^{2}}\right)^{3}\right)\ ,\end{split}$ (113)
$\frac{\nu_{2,\ell}}{\mathcal{T}}\bigg|_{d=4}$ $=\ell-\frac{6(\ell-1)}{\ell+1}\left(\frac{\lambda n}{16\pi^{2}}\right)$
$+\frac{3(\ell-1)(\ell(\ell(\ell(17\ell+78)+135)+98)+12)}{\ell(\ell+1)^{3}(\ell+2)}\left(\frac{\lambda n}{16\pi^{2}}\right)^{2}+\mathcal{O}\left(\left(\frac{\lambda n}{16\pi^{2}}\right)^{3}\right)\ .$ (114)
@@
$+\bigg[\frac{3n^{2}}{2(N+8)}-\bigg(\frac{4-N}{2(N+8)}+\sum_{\ell=1}^{\infty}\frac{q_{1,\ell}(1+3\ell)+3q_{2,\ell}(\ell-1)}{(1+\ell)(8+N)}\bigg)n+\mathcal{O}(n^{0})\bigg]\epsilon$
@@ 5.5.1 Ising CFT
$\Delta_{n}=n\left(1-\frac{\epsilon}{2}\right)+\frac{1}{6}(n-1)n\epsilon-\frac{n^{2}(17n-67)\epsilon^{2}}{324}+\mathcal{O}(\epsilon^{2}n,\epsilon^{3})\ ,$ (117)
which matches the diagrammatic result given in Eq. (A) with $N=1$ for the tower of operators of the form $\phi^{n}$. In the first row of Table 1, we display the semiclassical result for these anomalous dimensions $\gamma_{n,q_{\ell}}$, defined as in Eq. (20), together with the corresponding value of $q_{\ell}$.
@@
To build the spectrum of excited states, we note that the stability angles are ordered as $\nu_{2,\ell+1}>\nu_{2,\ell}$ and stress that $q_{2,\ell}$ for each $\ell$ has a given integer value. Intuitively, exciting an $\ell=\ell^{*}$ mode adds $\ell^{*}$ derivatives, in the $(\ell^{*}/2,\ell^{*}/2)$ representations of $SO(3,1)\cong SL(2,\mathbb{C})/Z_{2}$, to the operator. For example, if a given mode $\ell^{\ast}$ is excited twice and no other $\ell$ mode is excited we have $q_{2,\ell}=2\delta_{\ell,\ell^{\ast}}$ with $\delta_{\ell,\ell^{\ast}}$ the Kronecker delta. From the group theory point of view we have to take the tensor product of the corresponding $(\ell^{*}/2,\ell^{*}/2)$ representations with itself and perform the decomposition into irreducible representations to determine the representations of the resulting operators. As a consequence, all the so-obtained representations remain degenerate to the NLO order in the semiclassical expansion.
@@
    $-\left(67+\frac{3(s-1)\left(17s^{4}+78s^{3}+135s^{2}+98s+12\right)}{s(s+1)^{3}(s+2)}\right)n^{2}\Bigg]\epsilon^{2}+\mathcal{O}(\epsilon n^{0},\epsilon^{2}n,\epsilon^{3})\,.$ (118)
@@
    $\Delta_{n,\delta_{\ell,s}}=2+n\left(1-\frac{\epsilon}{2}\right)+\frac{(n-2)(3n+1)}{18}\epsilon-\frac{(n-2)(102n^{2}-335n-235)}{1944}\epsilon^{2}+\mathcal{O}(\epsilon^{3})\,.$ (119)
@@
  • $q_{2,\ell}=2\delta_{\ell,2}$ : Exciting an $\ell=2$ mode twice we have to take the tensor product of the two spin-2 representations (with highest weights $1$) $1\otimes 1=2\oplus 1\oplus 0$. Correspondingly, spin-4 $\partial^{4}\phi^{n}$, spin-2 $\partial^{2}\Box\phi^{n}$ and spin-0 $\Box^{2}\phi^{n}$ components will be degenerate to this semiclassical order with conformal dimension:
    $\Delta_{n,2\delta_{\ell,2}}=4+n\left(1-\frac{\epsilon}{2}\right)+\frac{n(3n-7)}{18}\epsilon-\frac{51n^{3}-338n}{972}\epsilon^{2}+\mathcal{O}(\epsilon n^{0},\epsilon^{2}n,\epsilon^{3})\,$, (120)
@@
    $\Delta_{n,\delta_{\ell,2}+\delta_{\ell,3}}=5+n\left(1-\frac{\epsilon}{2}\right)+\frac{n(3n-8)\epsilon}{18}-\frac{n^{2}(2040n-14533)\epsilon^{2}}{38880}+\mathcal{O}(\epsilon n^{0},\epsilon^{2}n,\epsilon^{3})\,$, (121)
@@
Evidently, by going further in our construction above one reproduces the full spectrum of primary operators in the traceless symmetric Lorentz representations whose spin is parametrically smaller than $n$. Remarkably, the framework naturally explains degeneracies in the Ising CFT spectrum Kehrein:1994ff ; Kehrein:1995ia which are harder to understand from a diagrammatic point of view. In particular, it is clear that the whole spectrum of operators becomes degenerate at the leading order in the semiclassical expansion, that is, in the $n\to\infty$ limit. In other words, perturbatively, the coefficient of the leading power of $n$ at any loop order is the same for all the families of operators and is equal to 1/6 and -17/324 at the one and two-loop orders, respectively (see Eq. (116)). The NLO semiclassical correction breaks the degeneracy among operators that can be distinguished by the choice of the $q_{\ell}$ integers. Remaining degeneracies are generically lifted by higher orders in the semiclassical expansion; see Table 2 for the NNLO $\mathcal{O}(n^{0})$ terms at $1$-loop order.
Finally, we remark that our results correspond to the eigenvalues of the Hamiltonian on the cylinder (i.e., energy spectrum) which via the state-operator correspondence are mapped to the eigenvalues of the dilatation operator on the plane (spectrum of anomalous dimensions). From the standard perturbative diagrammatic expansion point of view, before comparing to our results, we have to solve the operator mixing problem. For example, at spin-$6$ level and to $1$-loop in perturbation theory, for sufficiently large $n$, there are four towers of operators (shown in red color in Table 1), three of which mix through a degree-3 polynomial, and one which remains rational. To compare with our results, the general expressions for the eigenvalues of the $3\times 3$ mixing matrix have to be expanded at large $n$, and the leading $\mathcal{O}\left({n^{2}}\right)$ and subleading $\mathcal{O}\left({n}\right)$ terms will be reproduced by our computation. The reason for this expansion is again the fact that the full $1$-loop expression for the conformal dimensions of the spinning excited states has the leading $\mathcal{O}(n^{2})$, subleading $\mathcal{O}(n)$ and subsubleading power $\mathcal{O}(n^{0})$ term while the ground state operators $\phi^{n}$ and scalars $\Box^{2}\phi^{n}$ have only leading $n^{2}$ and subleading $n$ powers, see Table 2. As this subsubleading $\mathcal{O}(n^{0})$ term is NNLO in our semiclassical expansion, the full $1$-loop expression for the conformal dimension cannot be compared directly to our results [^6] We thank J. Henriksson for cross-checking the corresponding terms in perturbative $1$-loop results with our predictions in Antipin:2025ilv ..

### v2

@@ 5.3 Fluctuation operators
$\xi_{\ell,\pm}(z)=\frac{H(z\pm\alpha,\,|\,m)}{\Theta(z\,|\,m)}\,e^{\mp\,z\,Z(\alpha\,|\,m)}\ ,$ (80)
where $H,\Theta,Z$ respectively denote the Jacobi Eta, Theta, and Zeta functions reviewed in appendix B, and $\alpha$ solves
@@
$\xi_{\ell,\pm}(z)=\frac{H(z\pm\alpha_{+}\,|\,m)H(z\pm\alpha_{-}\,|\,m)}{\Theta(z\,|\,m)^{2}}e^{\mp z(Z(\alpha_{+}\,|\,m)+Z(\alpha_{-}\,|\,m))}\,$, (83)
@@
$\xi_{\ell,\pm}(z)=\prod_{j=1}^{\kappa}\frac{H(z\pm\alpha_{j}\,|\,m)}{\Theta(z\,|\,m)}e^{\mp zZ(\alpha_{j}\,|\,m)},$ (88)
@@
$C_{1}=r\delta E_{1}+\frac{r}{2{\cal T}}\sum_{\ell=0}^{\infty}\left(n_{\ell}\left[(N-1)\nu_{1,\ell}+\nu_{2,\ell}-N\nu_{0,\ell}\right]+2q_{1,\ell}\nu_{1,\ell}+2q_{2,\ell}\nu_{2,\ell}\right)\ ,$ (90)
@@
$C_{1}=r\delta E_{1}+\frac{r}{2{\cal T}}\sum_{\ell=0}^{\infty}\left(n_{\ell}\left[(N-1)\nu_{1,\ell}+\nu_{2,\ell}\right]+2q_{1,\ell}\nu_{1,\ell}+2q_{2,\ell}\nu_{2,\ell}\right)\ ,$ (94)
@@
We now comment on the zero modes of the theory. It is shown that when a symmetry of the theory is broken by a classical field configuration, a given pair of stability angles vanishes dhn . In our case, the stability angle $\nu_{2,0}$ corresponds to the zero mode related to the time translation symmetry on the cylinder stemming from the $t_{0}$ parameter in Eq. (9). The unnormalized eigenfunction is $\dot{v}(t)$ multiplied by the $\ell=0$ spherical harmonic. Indeed, one has $\Lambda_{2}(\ell=0)=1+4m$, which coincides with one of the band edges of the $L_{2}$ Lamé operator $L_{2}$. To evaluate the associated stability angle, we use that for $\ell=0$ and $d=4$ the equations (84) reduce to
@@ 5.4 Leading quantum correction: $C_{1}$
$\frac{1}{2}\sum_{\ell=0}^{\infty}n_{\ell}\nu_{2,\ell}=2{\cal T}/r+\frac{1}{2}\sum_{\ell=2}^{\infty}n_{\ell}\nu_{2,\ell}\ ,$ (101)
@@
$\frac{1}{2}\sum_{k=1}^{5}c_{k}\zeta(k-d)$ $=\pi-\frac{3((m-1)(3m-2)\mathbb{K}(m)+(4m-2){\mathbb{E}}(m))}{(1-2m)^{3/2}\epsilon}$
$-\frac{2(m-2)\mathbb{K}(m)+6{\mathbb{E}}(m)}{\sqrt{1-2m}}+\mathcal{O}(\epsilon)\ .$ (105)
@@
$-\frac{1}{2}\sum_{k=1}^{5}c_{k}|_{d=4}=\frac{(m(13-31m)+2)\mathbb{K}(m)+4\pi(1-2m)^{3/2}-(18-36m){\mathbb{E}}(m)}{(1-2m)^{3/2}}+\mathcal{O}(\epsilon)\ .$ (106)
@@
$-\frac{((29-5m)m-14)\mathbb{K}(m)-5\pi(1-2m)^{3/2}+(24-48m){\mathbb{E}}(m)}{(1-2m)^{3/2}}+\mathcal{O}(\epsilon)\ ,$ (107)
@@
By comparing the sum of Eq. (5.4) and Eq. (5.4) (multiplied by N1N-1) to Eq. (66), one can see that the divergent $1/\epsilon$ pole cancels in the renormalized combination $\mathcal{S}_{\rm cl}-\frac{1}{2}\sum_{\ell=0}^{\infty}n_{\ell}\left(\nu_{2,\ell}+(N-1)\nu_{1,\ell}\right)$. Dividing by the period $\cal T$}, adding our results for $\delta E_{1}$ in Eq. (69), and finally taking the $\epsilon\to 0$ limit, we arrive at
$C_{1}$ $=\frac{1}{8(1-2m)^{2}\mathbb{K}(m)}\Bigg[\mathbb{K}(m)(m(-7mN+26m+3N-70)+28)$
@@ 5.5 Perturbative semiclassics
$\frac{1}{2}\sum_{\ell=1}^{\infty}\sigma_{1}(\ell)$ $=\frac{11\pi m^{2}}{64}+\mathcal{O}\left(m^{3}\right)=11\pi\left(\frac{\lambda n}{16\pi^{2}}\right)^{2}+\mathcal{O}\left(\left(\frac{\lambda n}{16\pi^{2}}\right)^{3}\right)\,$, (112)
$\frac{1}{2}\sum_{\ell=2}^{\infty}\sigma_{2}(\ell)$ $=\frac{51\pi m^{2}}{64}+\mathcal{O}\left(m^{3}\right)=51\pi\left(\frac{\lambda n}{16\pi^{2}}\right)^{2}+\mathcal{O}\left(\left(\frac{\lambda n}{16\pi^{2}}\right)^{3}\right)\ ,$ (113)
@@
$\frac{r\nu_{1,\ell}}{\mathcal{T}}\bigg|_{d=4}$ $=\ell-\frac{2(3\ell+1)}{\ell+1}\left(\frac{\lambda n}{16\pi^{2}}\right)$
$-\left(\frac{20}{\ell+1}+\frac{8}{(\ell+1)^{3}}+\frac{2}{\ell+2}+\frac{2}{\ell}-51\right)\left(\frac{\lambda n}{16\pi^{2}}\right)^{2}+\mathcal{O}\left(\left(\frac{\lambda n}{16\pi^{2}}\right)^{3}\right)\ ,$ (114)
$\frac{r\nu_{2,\ell}}{\mathcal{T}}\bigg|_{d=4}$ $=\ell-\frac{6(\ell-1)}{\ell+1}\left(\frac{\lambda n}{16\pi^{2}}\right)$
$-\left(\frac{36}{\ell+1}+\frac{72}{(\ell+1)^{3}}+\frac{18}{\ell+2}+\frac{18}{\ell}-51\right)\left(\frac{\lambda n}{16\pi^{2}}\right)^{2}+\mathcal{O}\left(\left(\frac{\lambda n}{16\pi^{2}}\right)^{3}\right)\ .$ (115)
@@
$+\bigg[\frac{3n^{2}}{2(N+8)}-\bigg(\frac{4-N}{2(N+8)}+\sum_{\ell=1}^{\infty}\frac{q_{1,\ell}(1+3\ell)+3q_{2,\ell}(\ell-1)}{(1+\ell)(8+N)}\bigg)n+\mathcal{O}\left(n^{0}\right)\bigg]\epsilon$
@@ 5.5.1 Ising CFT
$\Delta_{n}=n\left(1-\frac{\epsilon}{2}\right)+\frac{1}{6}(n-1)n\epsilon-\frac{n^{2}(17n-67)\epsilon^{2}}{324}+\mathcal{O}\left(\epsilon^{2}n,\epsilon^{3}\right)\ ,$ (118)
which matches the diagrammatic result given in Eq. (A) with $N=1$ for the tower of operators of the form $\phi^{n}$. In the first row of Table 1, we display the semiclassical result for these anomalous dimensions $\gamma_{n,q_{\ell}}$, defined as in Eq. (21), together with the corresponding value of $q_{2,\ell}$.
@@
To build the spectrum of excited states, we note that the stability angles are ordered as $\nu_{2,\ell+1}>\nu_{2,\ell}$ and stress that $q_{2,\ell}$ for each $\ell$ has a given integer value. Intuitively, exciting an $\ell=\ell^{*}$ mode adds $\ell^{*}$ derivatives, in the $(\ell^{*}/2,\ell^{*}/2)$ representation of $SO(3,1)\cong SL(2,\mathbb{C})/Z_{2}$, to the operator. In particular, as evident from (117), for an operator of the form $\partial^{s}\Box^{p}\phi^{n}$ one has $\sum_{\ell=1}^{\infty}q_{2,\ell}\ell=2p+s$  [^6] For the $O(N)$ model this relation trivially generalizes as $\sum_{\ell=1}^{\infty}\left(q_{1,\ell}+q_{2,\ell}\right)\ell=2p+s$.. For example, if a given mode $\ell^{\ast}$ is excited twice and no other $\ell$ mode is excited we have $q_{2,\ell}=2\delta_{\ell,\ell^{\ast}}$ with $\delta_{\ell,\ell^{\ast}}$ the Kronecker delta. From the group theory point of view, we have to take the tensor product of the corresponding $(\ell^{*}/2,\ell^{*}/2)$ representations with itself and perform the decomposition into irreducible representations to determine the representations of the resulting operators. As a consequence, all the so-obtained representations remain degenerate to NLO in the semiclassical expansion.
@@
    $-\left(67+\frac{3(s-1)\left(17s^{4}+78s^{3}+135s^{2}+98s+12\right)}{s(s+1)^{3}(s+2)}\right)n^{2}\Bigg]\epsilon^{2}+\mathcal{O}\left(\epsilon n^{0},\epsilon^{2}n,\epsilon^{3}\right)\,.$ (119)
@@
    $\Delta_{n,\delta_{\ell,s}}=2+n\left(1-\frac{\epsilon}{2}\right)+\frac{(n-2)(3n+1)}{18}\epsilon-\frac{(n-2)(102n^{2}-335n-235)}{1944}\epsilon^{2}+\mathcal{O}\left(\epsilon^{3}\right)\,.$ (120)
@@
  • $q_{2,\ell}=2\delta_{\ell,2}$ : Exciting an $\ell=2$ mode twice we have to take the tensor product of the two spin-2 representations (with highest weights $1$) $1\otimes 1=2\oplus 1\oplus 0$. Correspondingly, the spin-4 $\partial^{4}\phi^{n}$, spin-2 $\partial^{2}\Box\phi^{n}$, and spin-0 $\Box^{2}\phi^{n}$ components will be degenerate to this semiclassical order with conformal dimension:
    $\Delta_{n,2\delta_{\ell,2}}=4+n\left(1-\frac{\epsilon}{2}\right)+\frac{n(3n-7)}{18}\epsilon-\frac{51n^{3}-338n}{972}\epsilon^{2}+\mathcal{O}\left(\epsilon n^{0},\epsilon^{2}n,\epsilon^{3}\right)\,$, (121)
@@
    $\Delta_{n,\delta_{\ell,2}+\delta_{\ell,3}}=5+n\left(1-\frac{\epsilon}{2}\right)+\frac{n(3n-8)\epsilon}{18}-\frac{n^{2}(2040n-14533)\epsilon^{2}}{38880}+\mathcal{O}\left(\epsilon n^{0},\epsilon^{2}n,\epsilon^{3}\right)\,$, (122)
@@
Evidently, by going further in our construction above one reproduces the full spectrum of primary operators in the traceless symmetric Lorentz representations whose spin is parametrically smaller than $n$. Remarkably, the framework naturally explains degeneracies in the Ising CFT spectrum Kehrein:1994ff ; Kehrein:1995ia which are harder to understand from a diagrammatic point of view. In particular, it is clear that the whole spectrum of operators becomes degenerate at the leading order in the semiclassical expansion, that is, in the $n\to\infty$ limit. In other words, perturbatively, the coefficient of the leading power of $n$ at any loop order is the same for all the families of operators and is equal to 1/6 and -17/324 at the one and two-loop orders, respectively (see Eq. (117)). The NLO semiclassical correction breaks the degeneracy among operators that can be distinguished by the choice of the $q_{\ell}$ integers. Remaining degeneracies are generically lifted by higher orders in the semiclassical expansion; see Table 2 for the NNLO $\mathcal{O}\left(n^{0}\right)$ terms at $1$-loop order.
Finally, we remark that our results correspond to the eigenvalues of the Hamiltonian on the cylinder (i.e., energy spectrum) which via the state-operator correspondence are mapped to the eigenvalues of the dilatation operator on the plane (spectrum of anomalous dimensions). From the standard perturbative diagrammatic expansion point of view, before comparing to our results, we have to solve the operator mixing problem. For example, at spin-$6$ level and to $1$-loop in perturbation theory, for sufficiently large $n$, there are four towers of operators (shown in red color in Table 1), three of which mix through a degree-3 polynomial, and one which remains rational. To compare with our results, the general expressions for the eigenvalues of the $3\times 3$ mixing matrix have to be expanded at large $n$, and the leading $\mathcal{O}\left({n^{2}}\right)$ and subleading $\mathcal{O}\left({n}\right)$ terms will be reproduced by our computation. The reason for this expansion is again the fact that the full $1$-loop expression for the conformal dimensions of the spinning excited states has the leading $\mathcal{O}\left(n^{2}\right)$, subleading $\mathcal{O}(n)$ and subsubleading power $\mathcal{O}\left(n^{0}\right)$ term while the ground state operators $\phi^{n}$ and scalars $\Box^{2}\phi^{n}$ have only leading $n^{2}$ and subleading $n$ powers, see Table 2. As this sub-sub-leading $\mathcal{O}\left(n^{0}\right)$ term is NNLO in our semiclassical expansion, the full $1$-loop expression for the conformal dimension cannot be compared directly to our results [^7] We thank J. Henriksson for cross-checking the corresponding terms in perturbative $1$-loop results with our predictions in Antipin:2025ilv ..